# TESSERA: Temporal Embeddings of Surface Spectra for Earth Representation and Analysis

### v1

@@ 2.6 Stocking indices in voluntary carbon markets
The most widely applied standard in the voluntary carbon market (VCM) for reforestation and afforestation projects of degraded pastures requires project proponents to demonstrate that their chosen remotely-sensed “stocking index” is correlated with the AGB (REF). Upon acceptance, project developers can use the index to quantify project additionality and to assess project performance in-between the 5-yearly ground-truth assessments. There is considerable freedom in choosing a stocking index, which can be either a pre-existing or proprietary canopy height or AGB product. Here, we evaluate TESSERA’s performance against five widely used, globally available canopy height and AGB products often used in the VCM. The in situ data used to compare the different stocking indices was collected by ICRAF in 38 sites in Para State, Brazil, that were degraded pastureland until being converted into agroforestry systems during 1980-2010 (Atzberger et al., 2025). To provide a fair comparison, stocking indices were rescaled using a linear transformation to minimize the RMSE with respect to the in situ data (details in the SI).
Figure 6: TESSERA outperforms commonly used stocking indices in the Voluntary Carbon Market. Comparison of in-situ aboveground biomass measurements in 38 agro-forestry plots in Para State, Brazil with TESSERA (top left) and five other remote sensing products. Note that the different products were originally provided in different units and converted to AGB as follows: (a) TESSERA: difference in relative heights of GEDI estimates of RH90 and RH10 converted to AGB using X * 15.502 + 160.5; (b) ETH canopy height converted using 3.4 X; (c) CTrees canopy height converted using 1.806 X + 44.9; (d) ESA: AGB converted using 0.407 X + 34; (e) Meta canopy height converted using 3.723 X + 46.6; (f) GEDI-Tandem-X above ground biomass prediction converted using 0.3X + 54.3. Details of conversation functions provided in SI
@@ 3 Conclusion
In this work, we introduce TESSERA, an innovative Remote Sensing Foundation Model that achieves worldwide coverage at 10-meter resolution through self-supervised learning applied to pixel-level satellite time series data. The representations generated by TESSERA establish new performance standards, while our open source methodology ensures widespread access to superior, high-resolution representations. We evaluated TESSERA’s capabilities across five distinct tasks, benchmarking our approach against leading task-specific models and existing foundation models. The findings demonstrate that TESSERA surpasses both conventional RF baselines and current state-of-the-art geospatial foundation models in these diverse downstream applications.
We next provide supplementary materials and methods about the details of how TESSERA is implemented.
Appendix A Earth Observation data
@@ A.1 Data Representation
We consider remote sensing data with $C$ channels (spectral bands or polarizations). Each data tile $R_{t}$ at time $t$ is represented as a 3D array with dimensions:
@@
where $V_{t}(i,j)=0$ indicates clouding or missing data for the pixel at spatial coordinates $(i,j)$, and $V_{t}(i,j)=1$ indicates valid data.
@@ A.2 Temporal Data Stacking
We stack spatially aligned tiles over a time period spanning $T$ time steps ($t=0,1,\ldots,T-1$). The temporal data stack is defined as:
@@ A.3 Time Series Extraction and d-pixel Definition
For a given spatial location $(i,j)$ and spectral channel $c$, the time series $S_{i,j,c}$ represents all channel $c$ values at coordinates $(i,j)$ over the entire time period:
@@
We define a d-pixel $P_{i,j}$ as the collection of all spectral channels by timesteps at a given spatial location $(i,j)$:
$P_{i,j}=S(i,j)$ (6)
In other words, the d-pixel provides all spectral values (Sentinel-2) or backscatter values (Sentinel-1) at a given point over time. Note that d-pixels are potentially sparse and have an accompanying mask $m_{i,j}$ that indicates the timesteps in which there is valid data:
$m_{i,j}=[v_{i,j,0},v_{i,j,1},\ldots,v_{i,j,T-1}]$ (7)
@@
$S_{i,j,c}$ $T\times 1$ Time series for channel $c$ at location $(i,j)$
$P_{i,j}$ $C\times T$ d-pixel: all spectral channels by timesteps at location $(i,j)$
$m_{i,j}$ $T\times 1$ d-pixel mask indicating valid timesteps at location $(i,j)$
@@
$C$ scalar Number of spectral channels
@@
Table 1: Variable definitions for remote sensing data structure
@@ A.4 The d-pixel Representation
We represent the rich information contained within satellite image time series at the level of individual pixels using the aforementioned “d-pixel”. For a single 10-meter geographic pixel location, it consolidates its complete annual multi-spectral (from Sentinel-2) and SAR backscatter (from Sentinel-1) time series. This structure explicitly preserves the inherent spectral and temporal dimensions of the data (Supplementary Fig. 1 a illustrates the d-pixel concept).
In this array, each row corresponds to a distinct observation date throughout the year, ordered chronologically. Each column corresponds to a specific spectral band or polarization. The value at the intersection of a given row and column thus represents the measured reflectance or backscatter for that specific band/polarization on that particular date for that single pixel.
A critical feature of the d-pixel is its capacity to accommodate the frequent data gaps encountered in remote sensing due to cloud cover (for optical data) or other atmospheric interference and acquisition irregularities. Observations identified as invalid (e.g., via Sentinel-2 scene classification layers for clouds, or missing Sentinel-1 acquisitions) are masked within the d-pixel array. These missing entries are explicitly handled by downstream model components, for instance, by being ignored during attention computation or managed during the temporal sampling process described in Appendix A.7.3.
@@ A.5 Dual-Encoder Architecture
  • Sentinel-2 MSI Encoder: This branch processes time series of 10 spectral bands from Sentinel-2. We used blue (B2), green (B3), red (B4), red edges 1–3 (B5, B6, B7), near-infrared (B8, B8A), and shortwave infrared (B11, B12).
  • Sentinel-1 SAR Encoder: This branch processes time series of 2 polarizations from Sentinel-1 (VV and VH).
Each encoder begins by linearly embedding the input features (spectral bands or polarizations) for each time step. To preserve sequence order and incorporate temporal context, learnable positional encodings based on the Day-of-Year (DOY) of each observation are added to these embeddings. The core of each encoder consists of a stack of 8 standard Transformer blocks [80], featuring multi-head self-attention and feed-forward layers to learn temporal patterns within the data streams.
To derive a single vector summarizing the entire time series for each modality, an attention-pooling layer weighs the importance of different time steps before aggregation. The resulting modality-specific representations (one from the S1 encoder, one from the S2 encoder) are then fused using a multi-layer perceptron.
@@ A.6 Projector Network
The fused representation from the dual-encoder stage is subsequently fed into a large projector network. This projector is a five-layer MLP, with each hidden layer having 16,384 dimensions. This significant expansion in dimensionality is crucial, as suggested by the original Barlow Twins work [90], to enable effective redundancy reduction during the self-supervised loss computation. The final output of the projector for each input d-pixel is an embedding, which is then used in the loss calculation. For downstream tasks, we typically use the 128-dimensional output from the fusion MLP (before the projector) as the final pixel representation. The TESSERA encoder (up to the fusion MLP) has approximately 40 million parameters, while the projector accounts for the majority of the model’s $\sim$1.4 billion parameters.
@@ A.7.2 Loss Function
Here, $C$ is the cross-correlation matrix computed between the batch-normalized embeddings $Z_{A}$ and $Z_{B}$. The indices $i$ and $j$ iterate over the dimensions of the embedding vectors. The first term (invariance term) encourages similar representations for different views of the same input ($C_{ii}\to 1$). The second term (redundancy reduction term) promotes informational efficiency by minimizing correlation between different embedding dimensions ($C_{ij}\to 0$ for $i\neq j$), weighted by $\lambda_{BT}$.
To further enhance model robustness and mitigate overfitting, TESSERA incorporates an additional mix-up regularization term, $\mathcal{L}_{MIX}$, inspired by Bandara et al. [9]. This involves shuffling one set of views (e.g., $Y_{B}$) along the batch dimension to create $Y_{S}=\text{Shuffle}(Y_{B})$, then generating mixed views $Y_{M}=\alpha_{mix}Y_{A}+(1-\alpha_{mix})Y_{S}$, where $\alpha_{mix}\sim\text{Beta}(\beta_{p},\beta_{p})$. The embeddings $Z_{M}$ and $Z_{S}$ are obtained. The mix-up loss penalizes deviations from the assumption that a linear interpolation in input space corresponds to a linear interpolation in embedding space:
@@
where $\lambda_{mix}$ controls the strength of the mix-up regularization. We found $\lambda_{BT}=5\times 10^{-3}$ and $\lambda_{mix}=1.0$ to be effective.
A.7.3 Pretraining Details
@@
The TESSERA model (Alpha version) was pretrained using approximately 0.8 billion d-pixel samples derived from 3,012 globally distributed Military Grid Reference System (MGRS) tiles. For pretraining, these d-pixels were generated from Sentinel-1 and Sentinel-2 data that were spatially downsampled by a factor of 400 from their native 10-meter resolution. Each d-pixel for Sentinel-2 contained time series for 10 spectral bands, and for Sentinel-1, 2 polarizations (VV and VH). The sequence length for the Transformer encoders was fixed at 40 timesteps for both modalities.
@@
For each pixel location, after filtering invalid observations (e.g., due to cloud cover for Sentinel-2), we performed sparse temporal sampling. This involves randomly selecting a fixed number of 40 valid observation dates from the year’s data. If fewer than 40 valid dates exist, sampling is performed with replacement. This strategy standardizes the input sequence length and serves as a key data augmentation mechanism, building invariance to data gaps and teaching the model that the underlying signal persists regardless of the specific dates observed. The temporal context of each sampled observation was encoded by transforming its normalized Day-of-Year (DOY) into sine and cosine features, which were then concatenated with the corresponding spectral or backscatter measurements. Finally, these values were standardized using global statistics to stabilize training.
@@
Figure 7: Geographical distribution and data density of the training dataset. TESSERA was trained on over 3,000 MGRS tiles distributed globally from 2017 to 2024. The color of each hexagon in the map corresponds to the number of valid observation days for Sentinel-1 (S1) and Sentinel-2 (S2), as defined by the bivariate color legend. This visualization highlights the density of combined S1 and S2 observations available for training across different regions, ensuring the model learns from a diverse range of geographical and environmental conditions.
@@
The model was trained for 1 epoch over the entire dataset, which corresponded to approximately 3,000 GPU hours on 8 AMD MI300X GPUs (192GB memory each). We used PyTorch with Fully Sharded Data Parallel (FSDP) and Automatic Mixed Precision (AMP) enabled. The AdamW optimizer was employed with a base learning rate of 0.002 and weight decay of $1\times 10^{-6}$. The learning rate schedule included a linear warmup over the initial 10% of steps, a plateau for the next 20%, followed by a cosine decay. The global batch size was 8192.
@@
A crucial aspect of our training methodology is a data shuffling strategy, essential for learning globally representative features from a vast and geographically diverse dataset. Given that d-pixels within an individual MGRS tile exhibit high spatial autocorrelation, a naive sequential or locally-shuffled data loading process would expose the model to strong geographic biases in each batch. To overcome this, we developed a custom data processing pipeline to implement a truly global shuffle across all $\approx$0.8 billion training samples, which constituted over 2TB of initial d-pixel data. This pipeline is illustrated conceptually in Fig. 8a.
@@
The impact of this approach on training stability is empirically demonstrated in Fig. 8b. Compared to a conventional, localized shuffling strategy which results in a highly volatile loss curve (top plot), our global shuffling strategy yields a markedly smoother and more stable convergence (bottom plot). This enhanced stability is fundamental for robust convergence and for preventing the model from overfitting to regional characteristics.
@@
Figure 8: Overview of the data shuffling pipeline and its impact on training stability. a, Schematic of the data shuffling and loading process. D-pixels (colored squares) from thousands of MGRS tiles are first aggregated into a global pool. A custom Rust binary performs a global shuffle on this multi-terabyte dataset before applying augmentations. The processed data is then organized into chunks and streamed by data workers to form well-shuffled, globally diverse training batches. b, Comparison of training loss curves. The top plot shows the volatile loss progression typical of a localized shuffling strategy, which is susceptible to geographic bias. The bottom plot shows the significantly smoother and more stable loss curve achieved with our global shuffling pipeline, demonstrating more effective and robust model convergence.
@@
To manage the significant I/O demands of shuffling and augmenting such a large volume of data, we developed the pipeline as a custom Rust binary. This high-performance binary handles the reading of raw d-pixel data, executes the global shuffle, and prepares the data for augmentation. The resulting pairs of augmented d-pixels are then serialized into a compact, pickle-like file format. These files are organized into manageable chunks and loaded by PyTorch DataLoader workers, which stream the data and assemble the final training batches. This end-to-end pipeline ensures that each batch presented to the model is a well-shuffled, globally diverse representation of Earth’s surface characteristics, which is fundamental for training a robust pixel-wise foundation model like TESSERA.
A.8 Global Representation Map Generation
A primary output of the TESSERA project is the generation of annual global representation maps at 10-meter resolution for the years 2017-2024.
A.8.1 Model Inference
To generate these maps, the pretrained and frozen TESSERA dual-encoder (excluding the projector) is used. For each 10-meter pixel on the globe and for each year:

### v2

@@ 2.5 Stocking indices in voluntary carbon markets
The conversion of degraded pastures into species-rich agroforestry systems has the potential to increase carbon sequestration while minimizing leakage liabilities [110]. To fund the transition, project developers can participate in the voluntary carbon market (VCM), but need to monitor above-ground biomass (AGB) over potentially large areas. This is a costly and inefficient undertaking if it is done only by field sampling.
To enable cost reductions, the most widely applied standard in the VCM (VM00047 from Verra [155]) allows project promoters to use remotely quantified ’stocking indices’ to assess project performance between 5-yearly ground truth assessments; project developers merely need to demonstrate that the chosen stocking index is correlated with the AGB. Upon acceptance, project developers can use the index to quantify the additionality of the project and to assess the performance of the project.
Here, we evaluate a simple TESSERA-derived stocking index against five widely used globally-available canopy height and AGB products that are often used in the VCM. To derive the stocking index from the TESSERA embeddings, GEDI relative height differences between RH95 and RH10 were used as target values for a pixel-wise RF regression following the methodology outlined in Atzberger et al. [13]. Without further fine-tuning, the derived stocking index, as well as five other state-of-the-art canopy height and AGB products (see  [1] for details), were individually rescaled so that their regression lines best matched (i.e., with slope 1 and zero bias) in situ data collected at 38 agroforestry sites in Para State, Brazil. This rescaling ensures a fair comparison between the various products.
A comparison of R2 and RMSE demonstrates that the TESSERA-derived stocking index consistently outperforms alternative products (Figure 6). As the alternative stocking indices all use similar (optical and microwave) inputs to derive canopy height, respectively, AGB, the superior performance of the TESSERA-derived stocking index must be attributed to the efficiency of the TESSERA embeddings.
Figure 6: TESSERA outperforms various stocking indices commonly used in the Voluntary Carbon Market. Comparison of in-situ AGB measurements in 38 agro-forestry plots in Para State, Brazil with TESSERA (top left) and five other remote sensing products (see  [1] for details). Note that the different products were originally provided in different units and converted to AGB as follows: (a) TESSERA: difference in relative heights of GEDI estimates of RH90 and RH10 converted to AGB using 15.502 X + 160.5; (b) ETH canopy height converted using 3.4 X; (c) CTrees canopy height converted using 1.806 X + 44.9; (d) ESA AGB converted using 0.407 X + 34.0; (e) Meta canopy height converted using 3.723 X + 46.6; (f) GEDI-Tandem-X AGB prediction converted using 0.3 X + 54.3. Details on the conversion functions are provided in Supplemental section I.
@@ 3 Conclusion
In this work, we introduce TESSERA, an innovative Remote Sensing Foundation Model that provides ready-to-use embeddings with worldwide coverage at 10-meter resolution through self-supervised learning applied to pixel-level optical and microwave satellite time series data. TESSERA embeddings establish new performance standards, while our open-source methodology and GeoTessera library permit use of our work even by non-specialists who lack access to high-performance computing. We evaluated TESSERA’s capabilities across five distinct tasks, benchmarking our approach against leading task-specific models and existing foundation models. Our findings demonstrate that TESSERA surpasses both conventional RF baselines and current state-of-the-art geospatial foundation models in these diverse downstream applications.
Acknowledgments
Funding:
ZF and SK were funded by a charitable donation made by Dr Robert Sansom to the University of Cambridge. SJ was funded by a charitable donation from John Bernstein. SS received funding from the Jenny and Antti Wihuri Foundation and the Helsinki Institute for Information Technology.
Author contributions:
TESSERA conceptualization: CA
System design: CA, ZF, SJ, SK
System implementation: ZF, RY, MCL, AM
Data curation: ZF, RY, JK, SJ, AM, DAC, SK
Downstream tasks: MCL, CA, SS, DAC, CA, MI, ZF, JK
Paper conceptualization: SK, AM, DAC, AB, SJ
Paper draft: ZF, MCL, RY, SS, SK, JK
Paper review and refinement: AB, CA, DAC, SJ, SK
Competing interests:
There are no competing interests to declare.
Data and materials availability:
Our framework integrates openly available data from the Copernicus Sentinel-2 (S2) [47] optical and Sentinel-1 (S1) [46] SAR missions, leveraging the strengths of each sensor. We used S2 Level-2A surface reflectance products, providing detailed spectral information across 10 bands (Red, Blue, Green, NIR, NIR8, Red-edge1, Red-edge2, Red-edge3, SWIR16, and SWIR22). To remove cloud contamination, we filtered the time series using the associated Sen2Cor scene classification map to mask cloudy observations. For SAR data, we processed Sentinel-1 Ground Range Detected imagery to derive calibrated backscatter coefficients in VV and VH polarizations, applying standard corrections using the ESA SNAP toolbox [48].
The Google Satellite Embedding V1 dataset [56] is licensed under CC-BY 4.0 and is produced by Google and Google DeepMind.
PRESTO embeddings can be created on demand in Google Earth Engine [57] using VertexAI [58].
Our global embedding map–initially for 2024 and eventually one for each year from 2017 to 2024–is $\sim$250 TB and is freely available for download, along with the parameters of the TESSERA model, the code to retrieve Sentinel-1 and Sentinel-2 data, and the training and inferencing code [152]. To allow for the seamless use of our embeddings in existing geospatial workflows, we have released the  geotessera Python library into the PyPI repository [124], which is the official third-party software repository for Python. This library automates the retrieval of TESSERA embeddings for a given year and geographical region, merging them into a numpy buffer [62] that can be passed directly into downstream tasks. The embeddings themselves are checksummed and indexed using Pooch [151] to check data integrity and handle retrieval.
The crop type identification downstream task uses the open INVEKOS Austrian crop dataset [3].
The burn area detection task uses the open Monitoring Trends in Burn Severity (MTBS) datasest [4].
The canopy height downstream task uses open access airborne laser scan data collected in the Danum valley and Sepilok [33].
The AGB downstream task uses the open BioMassters dataset [111].
The stocking index downstream task uses plot data from 38 plots in Brazil. This contains private information regarding plot owners, hence data access will require a Data Transfer Agreement.
Appendix A Model details
@@ A.1 Data Representation
We consider remote sensing data [46, 47] with $C$ channels (spectral bands or polarizations). Each data tile $R_{t}$ at time $t$ is represented as a 3D array with dimensions:
@@
[…]
@@ A.2 Temporal Data Stacking
We stack spatially aligned tiles over a time period that spans $T$ time steps ($t=0,1,\ldots,T-1$). The temporal data stack is defined as:
@@ A.3 Time Series Extraction and d-pixel Definition
[…]
@@
[…]
$P_{i,j}(c)=S(i,j,c)$ (6)
In other words, the d-pixel provides all spectral values (Sentinel-2) or backscatter values (Sentinel-1) at a given point over time. Note that d-pixels can be sparse and are accompanied by a mask vector $m_{i,j}$ of size $T$ that indicates the timesteps for which there is valid data, with a value 1 indicating that the corresponding row in $P_{i,j}$ is valid.
Table 1: Variable definitions for remote sensing data structure
@@
[…]
@@
$C$ Scalar Number of spectral channels + polarizations
@@ A.4 The d-pixel Representation
We represent each 10-m pixel in the time series of images from annual multi-spectral (Sentinel-2) and SAR backscatter (Sentinel-1) as a “d-pixel” array (see Fig. 1 a). In this array, each row corresponds to a chronologically ordered observation date, and each column is a specific spectral band or polarization. This representation accommodates data gaps encountered in remote sensing due to cloud cover (for optical data) or other atmospheric interference and acquisition irregularities. Observations identified as invalid (e.g., via Sentinel-2 scene classification layers for clouds, or missing Sentinel-1 acquisitions) are masked and are handled by downstream model components, for instance, by being ignored during attention computation or ignored in the temporal sampling process.
@@ A.5 Dual-Encoder Architecture
  • Sentinel-2 MSI Encoder: This branch processes a time series of 10 spectral bands (excluding the 60 m bands used for the detection of water vapor and clouds) from Sentinel-2. We used blue (B2), green (B3), red (B4), red edges 1–3 (B5, B6, B7), near-infrared (B8, B8A), and shortwave infrared (B11, B12).
  • Sentinel-1 SAR Encoder: This branch processes a time series of 2 polarizations from Sentinel-1 (VV and VH).
Each encoder begins by linearly embedding the input features (spectral bands or polarizations) for each time step. To preserve sequence order and incorporate temporal context, learnable positional encodings based on the Day-of-Year (DOY) of each observation are added to these embeddings. The core of each encoder consists of a stack of 4 standard Transformer blocks [153], featuring multi-head self-attention and feed-forward layers to learn temporal patterns within the data streams.
To derive a single vector summarizing the entire time series for each modality, an attention-pooling layer weighs the importance of different time steps before aggregation. The resulting modality-specific embeddings (one from the S1 encoder, one from the S2 encoder) are then fused using a multi-layer perceptron.
@@ A.6 Projector Network
The fused embedding from the dual-encoder stage is subsequently fed into a large projector network. This projector is a five-layer MLP, with each hidden layer having 16,384 dimensions. This significant expansion in dimensionality is crucial, as suggested by the original Barlow Twins work [187], to enable effective redundancy reduction during the self-supervised loss computation. The final output of the projector for each input d-pixel is an embedding, which is then used in the loss calculation. For downstream tasks, we use the 128-dimensional output from the fusion MLP (before the projector) as the final pixel embedding. The TESSERA encoder (up to the fusion MLP) has approximately 46 million parameters, while the projector accounts for the majority of the model’s $\sim$1.4 billion parameters.
Table 2: Architectural parameters of the TESSERA v1.0 model.
Component Parameters Percentage of Total
Encoder (Feature Extractor) 45,697,026 3.29%
   Sentinel-1 Encoder 22,846,465 1.64%
   Sentinel-2 Encoder 22,850,561 1.64%
Fusion MLP 131,200 $<$0.01%
Projector (for Training only) 1,344,700,421 96.71%
Total Parameters 1,390,528,647 100.00%
@@ A.7.2 Loss Function
Here, $C$ is the cross-correlation matrix computed between the batch-normalized embeddings $Z_{A}$ and $Z_{B}$. The indices $i$ and $j$ iterate over the dimensions of the embedding vectors. The first term (invariance term) encourages similar embeddings for different views of the same input ($C_{ii}\to 1$). The second term (redundancy reduction term) promotes information efficiency by minimizing the correlation between different embedding dimensions ($C_{ij}\to 0$ for $i\neq j$), weighted by $\lambda_{BT}$. We used the value of $\lambda_{BT}=5\times 10^{-3}$ as recommended in Reference [187].
To further enhance the robustness of the model and mitigate overfitting, TESSERA incorporates an additional mix-up regularization term $\mathcal{L}_{MIX}$, inspired by Bandara et al. [16]. This involves shuffling one set of views (e.g., $Y_{B}$) along the batch dimension to create $Y_{S}=\text{Shuffle}(Y_{B})$, then generating mixed views $Y_{M}=\alpha_{mix}Y_{A}+(1-\alpha_{mix})Y_{S}$, where $\alpha_{mix}\sim\text{Beta}(\beta_{p},\beta_{p})$. The embeddings $Z_{M}$ and $Z_{S}$ are obtained. The mix-up loss penalizes deviations from the assumption that a linear interpolation in the input space corresponds to a linear interpolation in the embedding space.
@@
where $\lambda_{mix}$ controls the strength of the mix-up regularization. We found $\lambda_{mix}=1.0$ (as recommended by [16]) to be effective based on our experiments.
A.7.3 Pre-training Details
@@
The TESSERA model was pre-trained using approximately 800 million d-pixels drawn from 3,012 globally distributed Military Grid Reference System (MGRS) tiles. d-pixels were generated by spatially downsampling Sentinel-1 and Sentinel-2 data by a factor of 400, i.e, selecting one pixel from each 20$\times$20 patch. Each d-pixel contained the annual time series for 10 spectral bands of Sentinel-2 data and the annual time series of VV and VH polarizations for Sentinel-1 data.
@@
For each pixel location, after removing invalid observations (e.g., due to cloud cover for Sentinel-2), we performed sparse temporal sampling. This involves randomly selecting a fixed number of 40 valid observation dates from the year’s data. The sequence length for the Transformer encoders was fixed at 40 timesteps for both modalities. If there are fewer than 40 cloud-free dates, replacement sampling is performed. This strategy standardizes the input sequence length and serves as a key data augmentation mechanism, building invariance to data gaps and teaching the model that the underlying signal persists regardless of the specific dates observed. Finally, these values were standardized using global statistics to stabilize training. The temporal context of each sampled observation was encoded by transforming its normalized Day-of-Year (DOY) into sine and cosine features, which were then concatenated with the corresponding spectral or backscatter measurements.
Table 3: Pre-training hyperparameters for the TESSERA v1.0 model.
Description Symbol Value
Optimizer & Schedule
Optimizer - AdamW
Batch Size (Per-GPU) $B_{local}$ 2,048
Batch Size (Global) $B_{global}$ 32,768 (16 GPUs $\times$ 2,048)
Epochs - 1
Base Learning Rate $\eta$ 0.002
Weight Decay $\lambda_{wd}$ $1\times 10^{-6}$
Learning Rate Schedule - Linear warmup (10%), plateau (20%), cosine decay
Gradient Clipping Norm - 2.0
Loss Function
Barlow Twins Lambda $\lambda_{BT}$ $5\times 10^{-3}$
Mixup Lambda $\lambda_{mix}$ 1.0
Mixup Beta Parameter $\beta_{p}$ 1.0
Architecture & Data
Sequence Length $L$ 40
Representation Dimension $D_{repr}$ 128
Projector Output Dimension $D_{proj}$ 16,384
Quantization Bits (QAT) - 8
@@
Figure 7: TESSERA is trained on globally-distributed training data with greater data density in the northern hemisphere. TESSERA was trained on over 3000 MGRS tiles distributed globally from 2017 to 2024. The color of each hexagon in the map corresponds to the number of valid observation days for Sentinel-1 (S1) and Sentinel-2 (S2), as defined by the bivariate color legend. This visualization highlights the density of combined S1 and S2 observations available for training across different regions, ensuring the model learns from a diverse range of geographical and environmental conditions.
@@
As we were in a compute-limited rather than data-limited pre-training regime, we trained for a single epoch, which corresponded to approximately 6200 GPU hours on 16 AMD MI300X GPUs (192GB memory each). We used PyTorch with Fully Sharded Data Parallel (FSDP) and Automatic Mixed Precision (AMP) enabled. AdamW optimizer was used with a base learning rate of 0.002 and weight decay of $1\times 10^{-6}$. The learning rate schedule included a linear warming during the initial 10% of steps, a plateau for the next 20%, followed by a cosine decay. The global batch size was 32768.
@@
A crucial aspect of our training methodology is a data shuffling strategy, essential for learning globally representative features from a representative sample of a vast and geographically diverse dataset. Given that d-pixels within an individual MGRS tile exhibit high spatial autocorrelation, a naive sequential or locally-shuffled data loading process would expose the model to strong geographic biases in each batch. To overcome this, we developed a custom data processing pipeline to implement a truly global shuffle across all $\approx$800 million training samples, which constituted more than 2TB of initial d-pixel data. This pipeline is conceptually illustrated in Fig. 8a.
@@
The impact of this approach on training stability is empirically demonstrated in Fig. 8b. Compared to a conventional, localized shuffling strategy, which results in a highly volatile loss curve (top plot), our global shuffling strategy yields a markedly smoother and more stable convergence (bottom plot). This enhanced stability is fundamental for robust convergence and for preventing the model from overfitting to regional characteristics.
@@
Figure 8: Data shuffling greatly improves training stability. a, Schematic of the data shuffling and loading process. D-pixels (colored squares) from thousands of MGRS tiles are first aggregated into a global pool. A custom Rust binary [152] performs a global shuffle on this multi-terabyte dataset before applying augmentations. The processed data is then organized into chunks and streamed by data workers to form well-shuffled, globally diverse training batches. b, Comparison of training loss curves. The top plot shows the volatile loss progression typical of a localized shuffling strategy, which is susceptible to geographic bias. The bottom plot shows the significantly smoother and more stable loss curve achieved with our global shuffling pipeline, demonstrating more effective and robust model convergence.
@@
To manage the significant I/O demands of shuffling and augmenting such a large volume of data, we developed a custom high-performance pipeline that handles the reading of raw d-pixel data, executes the global shuffle, and prepares the data for augmentation. The resulting pairs of augmented d-pixels are then serialized into a compact, pickle-like file format. These files are organized into manageable chunks and loaded by PyTorch DataLoader workers, which stream the data and assemble the final training batches. This end-to-end pipeline ensures that each batch presented to the model is a well-shuffled, globally diverse representation of Earth’s surface characteristics, which is fundamental for training a robust pixel-wise foundation model like TESSERA.
A.8 Global Embedding Map Generation
A primary output of the TESSERA project is the generation of annual global embedding maps with 10-meter resolution for the years 2017-2024. To generate these maps, the pre-trained and frozen TESSERA dual encoder (excluding the projector) is used. For each 10-meter pixel on the globe and for each year: